in endurance horse racing, people over the age of 15 are called seniors and people 15 years or younger are called juniors. An endurance horse race consists of 25 seniors and 5 juniors. (Give answer as a fraction or a decimal out to at least 4 places. If your answer is very small use scientific notation for example 3.3421E-6.)
endurance horse race
What is the probability that the top three finishers were:
(a) all seniors

(b) all juniors

(c) 2 seniors and 1 junior

(d) 1 senior and 2 juniors

Answer by Boreal(15235)

There are 30 in the race and 30C3 are the number of ways 3 finishers may occur. That is the denominator.
25C3 are the number of ways to choose 3 seniors from 25
so 25C3/30C3=0.5665
-
all juniors is 5C3/30C3=0.00246
-
2 seniors and 1 junior is 25C2*5C1/30C3=0.3695
-
1 senior and 2 juniors is is 25C1*5C230C3=250/30C3=0.06158